# Visual Planning: Let’s Think Only with Images

### v1

@@ 4 Discussions and Analysis
Error Analysis and Case Study. Figure 3 presents visual planning traces generated by LVM across different tasks. As defined in Section 2, the model occasionally takes non-optimal actions that deviate from the shortest path, as seen in the FrozenLake example. Invalid actions include violations of physical constraints (e.g., walking through walls in Maze or entering the table in MiniBehavior), or executing multiple actions in a single step (see Appendix C.5 for examples).
Figure 4 compares visual planning with language-based reasoning systems. In FrozenLake, Gemini 2.5 Pro misinterprets the environment size at the first step, causing cascading errors that lead to concluding an incorrect final answer. Similarly, the language-based SFT baseline makes an invalid action at the third step, reflecting difficulty in tracking states during reasoning. In contrast, visual planning avoids such failures by reasoning directly in the visual modality while reflecting the visual states per action. VPRL demonstrates the ability to take detours to bypass obstacles while still progressing toward the goal, whereas VPFT, lacking this flexibility, gets stuck and fails to reach the destination. More examples are provided in Appendix C.5.
@@
Random Policy Initialization Enables Exploration. A natural follow-up question arises: could we directly use VPFT as the policy model for GRPO training rather than intentionally initialize a model with random trajectories? We hypothesize that VPFT, trained via teacher-forcing, inherently limits exploration by repeatedly generating similar actions, resulting in identical rewards. In this case, it yields zero advantage, preventing policy updates and hindering effective learning. We empirically validate this hypothesis by comparing the exploration capabilities of VPFT with VPRL Stage 1 (Figure 6). We observe that VPFT’s entropy rapidly declines throughout training, eventually approaching zero, indicating severe exploration limitations. Although earlier VPFT checkpoints exhibit higher entropy, they produce significantly more invalid actions. In contrast, VPRL Stage 1 demonstrates significantly higher entropy, closely approaching the entropy of the uniform random planner, while maintaining a much lower invalid action ratio. These results justify the necessity of random initialization in our reinforcement learning framework to ensure robust exploration.
VPRL Reduces Invalid Action Failure. Another important benefit of VPRL lies in its effectiveness in reducing invalid actions. To quantify this, we analyze all failed trajectories and compute the proportion that contains at least one invalid action, as opposed to failures caused by non-optimal but valid plans. We refer to this as the invalid-failure ratio. As shown in Table 5, VPFT exhibits high ratio ranging from 61% to 78% over three tasks, while VPRL reduces this ratio by at least 24% in all cases, demonstrating that VPRL not only improves success rates, but also encourage the model to stay within valid action spaces during planning.
@@ 5 Related Work
Reinforcement learning has been applied across a wide range of vision-related tasks, especially given the rise of GRPO as in DeepSeek-R1 [16]. Concurrently, in object detection, visual perception [54] is optimized though rewarding high Intersection-over-Union (IoU) scores between predicted and ground-truth bounding boxes [45]. For visual reasoning tasks such as Visual Question Answering (VQA), GRPO has been utilized to optimize the models for longer, more coherent, and logically grounded reasoning traces in textual responses [34, 60, 58, 46]. More recently, similar methods have also been applied to image generation tasks, where the model is guided to reflect on the generated images and refine them recursively based on the alignment with given textual instructions [17, 48, 24]. These approaches focus on pixel-level fidelity and semantic alignment with text, whereas our work leverages RL for goal-oriented visual planning, optimizing multi-step decision-making through visual state transitions without any reliance on language. While prior RL-based approaches ground reasoning traces in textual outputs despite multimodal inputs, the modality mismatch limits the effectiveness of RL in bridging perception and action. For the tasks that are ‘vision-first’ by design, our visual planning paradigm and two-stage training framework VPRL enable more natural and flexible policy exploration by operating entirely in the visual domain, outperforming all language-based training variants.
@@ 6 Conclusion
In this work, we present Visual Planning as a new paradigm for reasoning in visually oriented tasks, challenging the prevailing reliance on language as the primary medium for structured inference. By enabling models to operate entirely through visual state transitions without textual mediation, we demonstrate that purely visual representations can lead to more effective and intuitive planning, particularly in spatially grounded and dynamic tasks. Our proposed two-stage reinforcement learning framework, VPRL, empowered by GRPO, further enhances the planning capabilities of large vision models. It obtains significant gains across three visual navigation tasks, achieving over 40% improvements in task performance than language-based planning and showing stronger generalization on out-of-distribution scenarios. These findings underscore the promise of visual planning as a powerful alternative to text-based approaches. We believe our work opens up a rich new direction for multimodal research, offering a foundation for building more intuitive, flexible, and powerful reasoning systems across a wide range of domains.
@@ Appendix A Limitations and future work
In this work, we focus exclusively on Large Vision Model (LVM) to investigate visual planning capabilities by eliminating language as a confounding factor for research purposes. As such, this choice constraints the model size to 3B as the only available size of LVM, and excludes recently released native multimodal models capable of generating multimodal outputs [9, 52]. However, we argue that the visual planning paradigm can be extended to broader multimodal generation models for use in more diverse tasks, combined with more modalities as long as they support image generation.
Additionally, explicitly generating images introduces computational overhead during inference compared to a textual response. However, we argue that language-based reasoning, especially for thinking models [13], can be equally or more time-consuming. In our demonstration, Gemini generated over 7,000 thinking tokens yet failed to provide the correct answer in the end. The computation overhead introduced by image generation can be alleviated through more compact image representations using fewer tokens [10], which we advocate for future research.
Another limitation in this work lies in the implementation of state-action parsing function. For simplicity, we adopt the rule-based approach that compares pixel-wise features between the current state and the previous state (details in Appendix B.3). While effective in our controlled setup, this method limits generalizability to broader task settings. Nevertheless, we argue that the core idea is extensible and could be supported by well-established computer vision techniques such as segmentation [41], contour detection [32] and etc. We encourage future research to explore more robust and scalable designs for visual state-action parsing to advance visual planning systems.
@@ B.1 Dataset
Dataset preparation. For both FrozenLake and Maze, we construct environments of grid sizes ranging from $3\times 3$ to $6\times 6$. For each size, we sample 1250 environments, with 1000 used for training and 250 held out for testing (Table 2). Each environment here is guaranteed to have a unique layout, and the agent is randomly initialized at a grid from which the goal is reachable, forming the initial state $v_{0}$. Due to the relatively limited diversity of environments layout in MiniBehavior, where the complexity arises primarily from the action space, sampling unique environments in a small grid size becomes challenging. Therefore we focus only on grid sizes $7\times 7$ and $8\times 8$, allowing duplicates in layout but varying agent spawn positions to ensure sufficient data volume. To prevent data leakage, we split the dataset based on layout identity, ensuring no layout overlap between the training and test sets.
@@
  • SFT in Text (Baseline): For each environment, we sample an optimal trajectory consisting a sequence of visual states $(v_{0},\ldots,v_{n})$ as the ground truth. Each transition between states is determined by an action, enabling us to derive a corresponding verbalied action sequence $(a_{0},\ldots,a_{n-1})$. The input to the model is formulated by concatenating a textual prompt with an image representation of the initial state $v_{0}$, while the target output is the verbalized action sequence representing the optimal trajectory. The detailed prompt is provided in Appendix D.
@@
Dataset Statistics. We evaluate the performance of different system variants with in-distribution and out-of-distribution (OOD) settings. Table 2 show the training data distribution over different grid sizes across three tasks. The numbers of training and testing samples for different system variants are shown in Table 3. For OOD evaluation, the enlarged grid sizes are shown in Table 7. OOD evaluation data includes 250 samples for each task.
B.2 Models
Large Vision Model (LVM) [4] is an autoregressive models for image generation, which is only pretrained with image sequences with no exposure to language data. The model uses a tokenizer based on the VQGAN architecture [12], which extracts visual information from raw images and encodes it into 256 tokens from a fixed codebook. The image is generated in an auto-regressive manner with discrete tokens, which are then fed into the image detokenizer. Although LVM supports multiple model sizes, only the 3B-parameter version is publicly available; thus, we use this variant in our experiments. For a fair comparison, we use Qwen 2.5-VL-Instruct [3] with a matching parameter size as our language-based baseline.
@@ B.3 Reward Implementation
We adopt rule-based state-action parsing function $\mathcal{P}$ and progress map $D(v)$ in VPRL. For progress map, we apply the Breadth First Search (BFS) to search for the optimal trajectories and calculates the progress at each position in the grid for each task. The progress map are then used as a reward signal to guide VPRL training.
Specifically, for state-action parsing function, we parse the state and identify the difference between current state and previous state through pixel-wise feature extractor. We first convert both input and predicted states into a coordinate-based representation by dividing the image into a grid based on its size. Each region corresponds to a discrete coordinate in the environment. To reduce sensitivity to color and focus on structural differences, we convert all images to grayscale. We subsequently compute the Intersection-over-Union (IoU) between each coordinate in the predicted state and the coordinate in the input state that contains the player (input coordinate). The coordinate in the predicted state with the highest IoU is selected as the predicted agent position. The action is then inferred by comparing the start and predicted positions according to task-specific movement rules. For example, in the Maze environment, movement across walls is not allowed and would be considered invalid.
@@
In MiniBehavior, we extend this logic to identify pick and drop actions. A pick is detected when the IoU between the printer’s location in the input and predicted states falls below a threshold, indicating that the printer has been removed. A drop is inferred when a coordinate corresponding to the table region shows a large MSE increase, suggesting the printer has been placed there. Additional edge cases in theses tasks are omitted for brevity.
For reward compuation, if the predicted action is valid, we compare the progress values from the progress map $D(v)$ between the input and predicted states. A reward of 1 is given if the predicted state shows greater progress toward the goal than the input state; otherwise, the reward is 0. Invalid actions are penalized with a reward of -5.
@@ B.4 Training details
Hyper-Parameters SFT in Text VPFT VPRL VPFT*
@@
Epochs 30 30 10 10 10 30
Learning Rate 1e-5 1.5e-4 1.5e-4 5e-5 1.5e-4 1.5e-4
Train Batch Size 16 8 8 1 8 8
Group Size N/A N/A N/A 10 N/A N/A
Grad Accumulation 2 1 1 1 1 1
GPUs 8 8 8 8 8 8
For all post-training experiments, we apply Low-Rank Adaptation (LoRA) [20] on both attention layers and feed-forward layers. The detailed hyper-parameters are shown in Table 4. Only the loss of the labels is calculated in an instruction-tuning manner [49] for SFT. The image tokenizer and detokenizer are frozen during training. We use the AdamW optimizer [35] for all training procedures.
When SFT for textual planning and visual planning, we train the model for a maximum of 30 epochs. For VPRL, we first do stage 1 on random trajectories for 10 epochs for the purpose of exploration. We then use GRPO to optimize the model for planning for another 10 epochs for stage 2. We sample a group of 10 candidate responses per prompt to compute the advantages accordingly. To encourage a balance between exploration and exploitation, we apply a KL divergence penalty with a coefficient $\beta=0.001$. We use TRL library for training [47]. We’ve conducted our experiments on the machine with 8$\times$×A100 GPUs.
@@ B.5 Licenses
Model-wise, Large Vision Model and Qwen 2.5 VL are under Apache-2.0 license. TRL is under Apache-2.0 license. We collect the Maze dataset with our own Python scripts. FrozenLake is collected from OpenAI Gym under the MIT License.
@@ C.1 Training
C.2 Performance with Scaling Difficulties

### v2

@@ 4 Discussions and Analysis
Error Analysis and Case Study. We conduct error analysis for language-based planning and visual planning. We observe that textual planning systems with both SFT and RL are prone to errors when grounding visual inputs to verbalized descriptions during the inference process, with 25.7% of generated coordinate-based layout descriptions and 22.3% of generated ASCII-based representations being mismatched with ground-truth layouts. Qualitative analysis of response from textual RL baselines (Figure 9 in Appendix C.2.1) and proprietary models (Figure 4) also reveal similar observations. Taken together, these results demonstrate an inherent modality gap where language may not be the most accurate and effective representation for vision-first problem. For visual planning, Figure 3 presents visual planning traces generated by LVM across different tasks. We observe that the model occasionally takes non-optimal actions that deviate from the shortest path (FrozenLake example). Surprisingly, VPRL demonstrates the ability to take detours to bypass the obstacles while still progressing towards the goal, whereas VPFT lacks this flexibility and gets stuck, as shown in Figure 4. Additional traces covering optimal, non-optimal, and invalid cases can be found in Appendix C.6. Beyond these in-domain analyses, we further evaluate generalization on larger unseen grids and perturbed image inputs, with results reported in Appendix C.4.
@@
Random policy initialization enables exploration. We ablate whether we could directly use VPFT as the policy model for GRPO training rather than intentionally initialize a model with random trajectories. We hypothesize that VPFT, trained via teacher-forcing, inherently limits exploration by repeatedly generating similar actions, resulting in identical rewards. In this case, it yields zero advantage, preventing policy updates and hindering effective learning. We empirically validate this hypothesis by comparing the exploration capabilities of VPFT with VPRL Stage 1 (Figure 6). We observe that VPFT’s entropy rapidly declines throughout training, eventually approaching zero, indicating severe exploration limitations. Although earlier VPFT checkpoints exhibit higher entropy, they produce significantly more invalid actions. In contrast, VPRL Stage 1 demonstrates significantly higher entropy, closely approaching the entropy of the uniform random planner, while maintaining a lower invalid action ratio, justifying the necessity of Stage 1 random initialization for RL framework.
VPRL reduces invalid action failure. Another important benefit of VPRL lies in its effectiveness in reducing invalid actions. To quantify this, we analyze all failed trajectories and compute the proportion that contains at least one invalid action, as opposed to failures caused by non-optimal but valid plans. We refer to this as the invalid-failure ratio. As shown in Table 6, VPFT exhibits a high ratio ranging from 61% to 78% over three tasks, while VPRL reduces this ratio by at least 24% in all cases, demonstrating that VPRL not only improves success rates, but also encourage the model to stay within valid action spaces during planning.
@@ 5 Related Work
Reinforcement learning has been applied across a wide range of vision-related tasks, especially given the rise of GRPO as in DeepSeek-R1 [17]. Concurrently, in object detection, visual perception [58] is optimized though rewarding high Intersection-over-Union (IoU) scores between predicted and ground-truth bounding boxes [49]. For visual reasoning tasks such as Visual Question Answering (VQA), GRPO has been utilized to optimize the models for longer, more coherent, and logically grounded reasoning traces in textual responses [37, 64, 62, 50]. More recently, similar methods have also been applied to image generation tasks for recursive refinement with textual instructions [18, 52, 27]. These approaches focus on pixel-level fidelity and semantic alignment with text, whereas our work leverages RL for goal-oriented visual planning, optimizing multi-step decision-making through visual state transitions without any reliance on language.
Action-conditional Generative Models
Action-conditional generative models has focused on constructing latent representations of the world and predicting future observations conditioned on given actions [20, 5]. These models learn transition dynamics and are central to model-based reinforcement learning, where they allow agents to simulate potential outcomes without interacting directly with the environment [21]. While effective for representation learning and short-horizon prediction, action-conditional generative models do not perform planning and must therefore be coupled with an external planner. In contrast, our approach constitutes a holistic planner that internalizes planning within the visual generative flow, which is more effective for visual tasks than traditional text-based planners that suffer from a modality gap. It can also benefit from action-conditional generative models by using predicted observations as inputs.
@@ 6 Conclusion
In this work, we present Visual Planning as a new paradigm for reasoning in visually oriented tasks, challenging the prevailing reliance on language as the primary medium for structured inference. By enabling models to operate entirely through visual state transitions without textual mediation, we show that purely visual representations provide performance comparable to text-based planning in spatially grounded and dynamic tasks, establishing visual planning as a viable alternative. More importantly, our proposed two-stage reinforcement learning framework, VPRL, empowered by GRPO, further enhances the planning capabilities of large vision models. It obtains significant gains across three visual navigation tasks, achieving 27% EM improvements in task performance than language-based planning and showing stronger generalization on out-of-distribution scenarios. These findings underscore the promise of visual planning as a powerful alternative to text-based approaches. We believe our work opens up a rich new direction for multimodal research, offering a foundation for building more intuitive, flexible, and powerful reasoning systems across a wide range of domains.
@@ Appendix A Limitations and future work
In this work, we focus exclusively on Large Vision Model (LVM) to investigate visual planning capabilities by eliminating language as a confounding factor for research purposes. As such, this choice constraints the model size to 7B as the only available size of LVM, and excludes recently released native multimodal models capable of generating multimodal outputs [10, 56]. However, we argue that the visual planning paradigm can be extended to broader multimodal generation models for use in more diverse tasks, combined with more modalities, as long as they support image generation.
Additionally, explicitly generating images introduces computational overhead during inference compared to a textual response. However, we argue that language-based reasoning performs worse than visual planning and can be equally or even more time-consuming, especially for thinking models [14]. In our demonstration, Gemini generated over 7,000 thinking tokens yet failed to provide the correct answer in the end. The computation overhead introduced by image generation can be alleviated through more compact image representations using fewer tokens [11], which we advocate for future research.
Another limitation in this work lies in the implementation of dynamics interpreter. For simplicity, we adopt the rule-based approach that compares pixel-wise features between the current state and the previous state (details in Appendix B.3). While effective in our controlled setup, broader task settings with more diverse types of dynamic interpreters are yet to be explored. Nevertheless, we argue that the core idea is extensible and could be supported by either dynamic models to elicit actions from pairs of images (e.g. segmentation [45] or contour detection [35]) or a holistic neural model (e.g. Gemini [14]) that directly judges whether the transitions are valid without explicitly inferring actions. We encourage future research to explore more robust and scalable designs for the visual dynamics interpreter to advance visual planning systems.
@@ B.1 Dataset
Dataset preparation. For both FrozenLake and Maze, we construct environments of grid sizes ranging from $3\times 3$ to $6\times 6$. For each size, we sample 1250 environments, with 1000 used for training and 250 held out for testing (Table 3). Each environment here is guaranteed to have a unique layout, and the agent is randomly initialized at a grid from which the goal is reachable, forming the initial state $v_{0}$. Due to the relatively limited diversity of environments layout in MiniBehavior, where the complexity arises primarily from the action space, sampling unique environments in a small grid size becomes challenging. Therefore, we focus only on grid sizes $7\times 7$ and $8\times 8$, allowing duplicates in layout but varying agent spawn positions to ensure sufficient data volume. To prevent data leakage, we split the dataset based on layout identity, ensuring no layout overlap between the training and test sets.
@@
  • SFT in Text (Baseline): For each environment, we sample an optimal trajectory consisting of a sequence of visual states $(v_{0},\ldots,v_{n})$ as the ground truth. Each transition between states is determined by an action, enabling us to derive a corresponding verbalized action sequence $(a_{0},\ldots,a_{n-1})$. The input to the model is formulated by concatenating a textual prompt with an image representation of the initial state $v_{0}$, while the target output is the verbalized action sequence representing the optimal trajectory. We further ablate different variants of the baseline with various representations and tuning methods (SFT and RL) in Appendix C.2. The detailed prompts for all variants are provided in Appendix D.
@@
Dataset Statistics.
@@
We evaluate the performance of different system variants in in-distribution and out-of-distribution (OOD) settings. Table 3 shows the training data distribution over different grid sizes across three tasks. The numbers of training and testing samples for different system variants are shown in Table 4. For OOD evaluation, the enlarged grid sizes are shown in Table 9. OOD evaluation data includes 250 samples for each task.
B.2 Models
Large Vision Model (LVM) [4] is an autoregressive models for image generation, which is only pretrained with image sequences with no exposure to language data. The model uses a tokenizer based on the VQGAN architecture [13], which extracts visual information from raw images and encodes it into 256 tokens from a fixed codebook. The image is generated in an auto-regressive manner with discrete tokens, which are then fed into the image detokenizer. Although LVM supports multiple model sizes, only the 7B-parameter version is publicly available; thus, we use this variant in our experiments. For a fair comparison, we use Qwen 2.5-VL-Instruct [3] with a matching parameter size as our language-based baseline.
@@ B.3 Reward Implementation
We adopt a rule-based state-action parsing function as the dynamics interpreter $\mathcal{D}$ and heuristic progress estimator $P$ in VPRL. For the progress estimator, we apply the Breadth First Search (BFS) to search for the optimal trajectories and calculate the progress at each position in the grid for each task, in order to obtain a progress map covering all positions. The progress map are then used as a reward signal to guide VPRL training.
Specifically, for state-action parsing function, we parse the state and identify the difference between the current state and the previous state through a pixel-wise feature extractor. We first convert both input and predicted states into a coordinate-based representation by dividing the image into a grid based on its size. Each region corresponds to a discrete coordinate in the environment. To reduce sensitivity to color and focus on structural differences, we convert all images to grayscale. We subsequently compute the Intersection-over-Union (IoU) between each coordinate in the predicted state and the coordinate in the input state that contains the player (input coordinate). The coordinate in the predicted state with the highest IoU is selected as the predicted agent position. The action is then inferred by comparing the start and predicted positions according to task-specific movement rules. For example, in the Maze environment, movement across walls is not allowed and would be considered invalid.
@@
In MiniBehavior, we extend this logic to identify pick and drop actions. A pick is detected when the IoU between the printer’s location in the input and predicted states falls below a threshold, indicating that the printer has been removed. A drop is inferred when a coordinate corresponding to the table region shows a large MSE increase, suggesting the printer has been placed there. Additional edge cases in these tasks are omitted for brevity.
For reward computation, if the predicted action is valid, we compare the progress values from the heuristic progress estimator $P$ between the input and predicted states. A reward of 1 is given if the predicted state shows greater progress toward the goal than the input state; otherwise, the reward is 0. Invalid actions are penalized with a reward of -5.
@@ B.4 Training details
Hyper-Parameters SFT in Text RL in Text VPFT VPRL VPFT*
@@
Epochs 30 10 30 10 10 10 30
Learning Rate 1e-5 1e-5 1.5e-4 1.5e-4 5e-5 1.5e-4 1.5e-4
Train Batch Size 4 1 8 8 1 8 8
Group Size N/A 8 N/A N/A 10 N/A N/A
Grad Accumulation 1 1 1 1 1 1 1
GPUs 8 8 8 8 8 8 8
Beyond directly training on action labels, we further conduct an ablation on FrozenLake with textual variants that verbalize the input state before predicting the action sequence. In particular, we explore two structured representations: Coordinate descriptions and ASCII grids. During training, the target sequence consists of the description tokens (encoding the environment layout in either coordinate or ASCII form) concatenated with the action labels that lead to the goal.
Reinforcement Learning (RL) in Text. We also extend RL to textual planning in the FrozenLake environment as an ablation. We optimize the textual planner with Group Relative Policy Optimization (GRPO). The reward design combines a fixed format reward, which enforces the correct output structure, with an outcome reward defined in two variants: (1) a progress-based reward identical to that used in VPRL, or (2) the Progress Rate (PR) metric directly used as the reward.
For all post-training experiments, we apply Low-Rank Adaptation (LoRA) [23] on both attention layers and feed-forward layers. The detailed hyper-parameters are shown in Table 5. Only the loss of the targets is calculated in an instruction-tuning manner [53] for SFT. The image tokenizer and detokenizer are frozen during training. We use the AdamW optimizer [38] for all training procedures.
[…]
@@ B.5 Licenses
Model-wise, Large Vision Model and Qwen 2.5 VL are under the Apache-2.0 license. TRL is under the Apache-2.0 license. We collect the Maze dataset with our own Python scripts. FrozenLake is collected from OpenAI Gym under the MIT License.
@@ C.1 VPRL Training
C.2 Trained textual baselines and reward design
Table 7: Performance of text-based variants of Qwen-2.5-VL-Instruct-3B and 7B on FrozenLake. We report Exact Match (EM) and Progress Rate (PR) across all difficulty levels (L3–L6) and their average.
Model / Variant EM (%) PR (%) L3 L4 L5 L6 Avg. L3 L4 L5 L6 Avg. Qwen 2.5-VL-Instruct-3B   xxx- SFT   xxxxxx- Direct 87.2 72.0 48.8 28.0 59.0 89.4 84.3 71.3 60.1 76.3   xxxxxx- w/ Coordinates 87.2 78.0 64.8 30.8 65.2 89.6 82.5 74.0 57.2 75.8   xxxxxx- w/ ASCII 79.6 75.6 58.8 34.0 62.0 83.3 82.5 74.8 59.1 74.9   xxx- GRPO   xxxxxx- w/ VPRL progress reward 69.2 52.8 41.2 26.0 47.3 73.6 72.2 66.2 55.0 66.8   xxxxxx- w/ PR metric reward 70.8 60.0 41.6 23.6 49.0 75.5 76.1 65.7 56.0 68.4 Qwen 2.5-VL-Instruct-7B   xxx- SFT   xxxxxx- Direct 97.6 86.0 56.4 34.4 68.6 98.1 92.1 78.9 68.4 84.4   xxxxxx- w/ Coordinates 93.2 88.0 74.8 41.6 74.4 94.1 89.7 81.5 65.5 82.7   xxxxxx- w/ ASCII 93.2 86.0 68.0 45.2 73.1 94.1 88.6 81.3 69.6 83.4   xxx- GRPO   xxxxxx- w/ VPRL progress reward 72.4 64.0 50.4 30.8 54.4 76.2 76.3 69.2 57.8 69.9   xxxxxx- w/ PR metric reward 82.8 68.8 51.6 37.2 60.1 84.9 79.6 71.5 61.0 74.3 LVM-7B   xxx- VPFT (ours) 92.0 82.8 68.8 58.0 75.4 93.1 84.7 73.4 66.9 79.5   xxx- VPRL (ours) 97.6 95.6 90.8 82.4 91.6 98.4 96.0 93.0 85.6 93.2
Exact Match (%)
Model $\times3$ $\times4$ $\times5$ $\times6$
VPFT* 86.4 73.6 50.0 33.2
VPFT 92.0 82.8 68.8 58.0
Model FrozenLake ($\times7$) Maze ($\times7$) MiniBehavior ($\times9$) EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) VPFT 9.6 15.3 9.2 17.8 0.0 5.8 VPRL 20.4 31.2 10.0 21.6 0.4 14.7
To strengthen the comparison with our visual planners, we train different text-based baselines beyond the direct action-sequence SFT model reported in the main paper. We are interested in: 1) whether different textual representation influences the performance of language-based reasoning, and 2) whether reinforcement learning can help to improve the language-based planning performance with multimodal input.
Trained SFT variants. Specifically, we experiment with two alternative SFT variants that first describe the environment layout in different formats (coordinates and ASCII) before predicting the action sequence.
  • SFT with Coordinates: The model is trained to first output a coordinate-based description of the grid environment (e.g., positions of the agent, goal, and obstacles), followed by the full action sequence.
  • SFT with ASCII: The model is trained to output an ASCII-based description of the environment layout before producing the action sequence. Specifically, S denotes the starting position, G the goal, H an ice hole, and F a passable cell.
The example input-output formats for different text-based reasoning variants are shown in Figure 8 in Appendix C.2.1.
We experiment with both variants for Qwen-2.5-VL-Instruct-3B and 7B, training them with the same configurations as the original text SFT baseline. As shown in Table 7, the SFT variants with either coordinates or ASCII do not provide consistent significant improvements over the direct SFT baseline. Specifically, these variants with additional structural descriptions in either coordinates or ASCII yield slight gains in EM, but exhibit lower PR compared to the direct SFT baseline. Moreover, both variants still fall short of VPRL, suggesting that enriching textual input alone is insufficient to bridge the gap between visual and text-based planning.
RL-trained text baseline. We also explore the feasibility of applying RL to improve the planning performance with multimodal input, given the success of RL in the pure language planning domain [17]. We train an RL-based text model using Qwen-2.5-VL-Instruct-3B and 7B with the GRPO algorithm, with output format shown in Figure 8.
We adopt the same progress-based reward design as in VPRL for fair comparison, in addition to a simple format reward that ensures reasoning is enclosed within <think> tag and the final answer within <answer> tag.
  • If the action is optimal (i.e., aligned with some optimal trajectory from the current state), it receives a reward of +1+1.
  • If the action is valid but non-optimal, it receives 0.
  • If the action is invalid (e.g., colliding with walls or stepping into holes), it receives 5-5, and all subsequent actions are also treated as invalid.
Input Example Refer to caption Task: Generate the shortest valid sequence of actions that moves the elf from the starting position to the goal without stepping into any ice holes (see detailed prompts in Appendix D).
SFT (w/ Coordinates) After analysis, the layout of the given image is as follows:
- Starting position: (3, 3)
- Goal: (2, 2)
- Ice holes: (1, 1), (2, 3)
- Passable cells: (1, 2), (1, 3), (2, 1), (3, 1), (3, 2)
Based on the above layout, the final answer is: <ANSWER>left up</ANSWER>
SFT (w/ ASCII) The symbols used to represent the grid are:
- S denotes the starting position, G the goal, H the ice hole, F passable cells.
HFH
FGH
FFS
Based on the above layout, the final answer is: <ANSWER>left up</ANSWER>
GRPO <think>
1. The elf is in the bottom right cell.
2. The goal is in the top center cell.
3. The top left cell has an ice hole, so the elf cannot go left there.
4. The elf needs to move one cell left to the center cell.
5. Then the elf needs to move one cell up to the goal cell.
The sequence of actions are:
1. Move one cell left to the center cell.
2. Move one cell up to the goal cell.
</think>
<answer>left up</answer>
Figure 8: Examples of model outputs under different trained textual variants: SFT (w/ Coordinate), SFT (w/ ASCII), and GRPO. Each variant follows a distinct format for representing the environment. Even for the relatively simple 3 × 3 example input, we observe that the ASCII and GRPO models still produce erroneous layout descriptions.
Level 3 Refer to caption <think>
1. The elf is at the center cell.
2. The goal is the gift box in the top left corner.
3. The elf needs to move 1 cell left and 1 cell up.
4. There are ice holes directly below and to the right, so they are impassable.
</think>
<answer>left up</answer>
Level 4 Refer to caption <think>
1. The elf is at the top left section of the grid.
2. The goal is the gift box in the bottom right section.
3. The elf needs to move 3 cells down and 1 cell right.
4. There are ice holes directly below and to the right, so the elf cannot move through them.
</think>
<answer>down down down right</answer>
Level 5 Refer to caption <think>
1. The elf is at the bottom left.
2. The goal is the gift box in the top right.
3. The elf needs to move 3 cells right and 2 cells up.
4. There are ice holes directly above and to the right, so the elf cannot move through them.
</think>
<answer>right right right up up</answer>
Level 6 Refer to caption <think>
1. The elf is at the top right.
2. The goal is the gift box at the bottom left.
3. The elf needs to move 3 cells down and 1 cell left.
4. There are ice holes directly below and to the right of the elf, so we cannot use those paths.
</think>
<answer>left down down down</answer>
Figure 9: Examples of responses from the textual GRPO baseline with PR metric as the reward on FrozenLake across different difficulty levels. Each box shows the input image and the corresponding model output. In all cases, the model produces incorrect layout descriptions, which in turn lead to incorrect predicted action sequences.
To prevent reward hacking through excessively short rollouts, any sequence shorter than the length of the optimal trajectory is padded with invalid actions for the remaining steps, which are considered “stay-in-place" moves, in other words, invalid transitions. Finally, to make rewards comparable across sequences of varying lengths, we normalize the total reward by sequence length $\sum_{t=1}^{T}r_{t}/T$.
We train the text RL baselines for 10 epochs, consistent with VPRL. Using the VPRL progress reward described above, the model achieves 54.4% EM (Table 7). We suspect that this limited performance is due to the reward design not being sufficiently discriminative. In particular, trajectories that contain the same number of optimal and non-optimal actions receive identical rewards regardless of their order (e.g., starting with optimal actions and then switching to non-optimal ones yields the same return as the reverse). As a result, the model tends to first learn to produce valid actions in general, and only later to distinguish optimal actions among them.
To address this issue, we further design an alternative reward function by directly adopting the Progress Rate (PR) metric from the main paper. This formulation encourages the model to focus on generating consecutive valid forward moves from the start, rather than separating the learning of validity and optimality. Under the same training conditions, this reward improves EM to 60.1%, but the performance still lags behind the direct SFT baseline. As we discussed in Section 4 (error analysis paragraph), we attribute the bottleneck of language-based planning with RL to the modality gap, which introduces inaccuracies in grounding visual information into text, causing exploration to proceed from misinterpreted states and thereby reducing the overall effectiveness of learning. By contrast, our visual planning paradigm avoids this modality gap by operating directly in the visual domain, ensuring exploration within the correct state space via RL.
C.2.1 Examples of trained Textual variants
Outputs of different textual variants are illustrated in Figure 8, including SFT with coordinate and ASCII representations, as well as GRPO with reasoning traces. Even for the relatively simple $3\times 3$ input, and despite all variants producing the correct final predictions shown in the figure, we observe that the ASCII and GRPO models still generate erroneous layout descriptions: in the ASCII case, the passable cell at the top right is misclassified as an ice hole, while in the GRPO case, the goal position is incorrectly identified.
We also conduct further qualitative analysis of responses from the textual RL baseline trained with the PR metric as the reward (Figure 9). In all cases, the model produces incorrect layout descriptions, which in turn lead to incorrect predicted action sequences, highlighting the modality gap in grounding visual information into text.
C.3 Performance with Scaling Difficulties